Hydrogen, iodine, and hydrogen iodide gases have reached equilibrium in a container at 700K. Equilibrium partial pressures are 0.28 atm HI, 0.033 atm H2, and 0.045 atm I2. For H2 (g) + I2 (g) --> 2HI, K=53. Half the HI is removed form the container and equilibrium is reestablished. Calculate the new partial pressure (in atm) of HI in the new equilibrium mixture.
1
Expert's answer
2012-11-27T08:03:21-0500
K = [HI]2/[I2][H2] = P2(HI)/P(I2)*P(H2) = 53
half of HI is proportional to pressure = 0.28/2 = 0.14

53 = (0.14+x2 )2 / (0.033-x)*(0.045-x)
x=....

new p of HI = 0.14+x (from upper equation)